Creativeone24:

Factor. x2−6x+9 (x+9)2 (x−9)2 (x+3)2 (x−3)2

Shadow:

Please do not give Direct Answers. Explain how you got there.

Creativeone24:

dnt worry i didnt even see it

Shadow:

@Creativeone24 Do you know how to factor?

Shadow:

For example, can you factor: \[x^2 + x\]

Creativeone24:

to ,figure out the factor i think ill have to factorize the quad in the equation.

Creativeone24:

oh wait it doesnt have a quad i messed up

Shadow:

Let me show you some examples

Creativeone24:

ok

Shadow:

\[x^2 + 6x + 8\] We need to do something called FOIL. Is this term familiar to you?

Creativeone24:

yes it is

Shadow:

F = First O = Outer I = Inner L = Last Say we have (a + b)(c + d) F means you multiply the first terms of each parentheses: a times c O means you multiply the outer term of each parentheses: a times d I means you multiply the inner terms of each parentheses: b times c L means you multiply the last terms of each parentheses: b times d

Shadow:

We want a setup that gives us: \[x^2 + 6x + 8\] \[x \times x = x^2 \] \[(x + ?)(x + ?)\] Whenever you see x^2, the first term of each parentheses is x. Now. The last two terms of each parentheses is always a factor of the last term of the polynomial. This is because you get the last term of the polynomial by multiplying the last terms of each parentheses ? times ? = 8 Factors of 8: 1,2,4,8 Now, which of these terms add up to 6. We can use this approach because the outer and inner terms will just be multiply by 1x, therefore they only get a variable and the coefficient does not change. 2 + 4 = 6 \[(x + 2)(x + 4)\]

Shadow:

Lets apply this same approach to: \[x^2−6x+9 \] We know we start with: \[(x + ?)(x + ?)\] Since we have x^2 Next, what is the last term of our polynomial? 9 Factors of 9: 1,3,9 Keep in mind that this always includes positive and negative values of the factors. So, -1, 1, -3, 3, -9, 9 So which of these factors add up to -6, and multiply to get positive 9? Why -3 and -3. -3 + -3 = -6 -3 times -3 = 9 neg times neg = pos Therefore, we get \[(x - 3)(x - 3)\]

Shadow:

Whenever you have this, it can be rewritten as: \[(x - 3)(x - 3) = (x -3)^2\]

Shadow:

@Creativeone24

Shadow:

Grabbing lunch, I hope this helps :)

Creativeone24:

thnk very much it does!!!

Creativeone24:

@Shadow